# Breaking Data Silos: Towards Open and Scalable Mobility Foundation Models via Generative Continual Learning

### v4

@@ 1. Introduction
In natural language processing (NLP) (Kojima et al., 2022; Brown et al., 2020; Bi et al., 2024) and computer vision (CV) (Esser et al., 2024; Liu et al., 2024b), the rise of large pre-trained foundation models (e.g., GPT, Sora) (Radford et al., 2021) has significantly advanced model sharing and general-purpose intelligence through centralized training and open release. However, in the critical domain of human mobility modeling (Barbosa et al., 2018; Luca et al., 2021), the era of large foundation models has not yet arrived. This is primarily due to the highly sensitive nature of mobility trajectory data, which involves personal privacy (Kong et al., 2023; Yuan et al., 2025a, 2022). Such data cannot be easily shared or jointly trained across institutions due to privacy constraints and legal restrictions. As a result, mobility datasets often exist in isolated silos (Yuan et al., 2025b; Zhu et al., 2023), with different organizations and researchers relying on their own private datasets, hindering data integration, benchmarking, and collaborative model development.
@@
To address the challenge of data silos in mobility modeling, recent studies have proposed different training strategies to leverage multiple datasets, and Figure 1 compares different approaches. The most common strategy is to train models separately, as shown in Figure 1(a). Recent efforts such as UniTraj (Zhu et al., 2024), TrajBert (Si et al., 2023), and TrajFM (Lin et al., 2024) have explored joint training for unified representation and cross-city generalization (Figure 1(b)), but these models remain tightly coupled with proprietary or limited-quality datasets. TrajFM and TrajBert are typically pre-trained on restricted or private data. While UniTraj uses public data, it suffers from low semantic richness and high sparsity. Federated learning (Feng et al., 2020; Liu et al., 2024a) offers a potential solution for distributed mobility model training (Figure 1(c)), but its reliance on frequent synchronization and communication poses challenges for scalability and practical deployment. Consequently, these approaches fall short of meeting the diverse, dynamic, and multi-source demands of real-world mobility modeling, and cannot support an open ecosystem of shared models as seen in NLP and CV domains.
@@
To address these challenges, we propose MoveGCL, a scalable training framework for mobility foundation models based on generative continual learning. MoveGCL allows each data holder to evolve a shared model locally without exposing raw data, thereby ensuring full privacy preservation. Specifically, MoveGCL starts from a pre-trained base model and employs a synthetic trajectory replay mechanism: instead of accessing historical data, each participant generates synthetic trajectories that approximate previously seen mobility patterns. This replay process preserves prior knowledge and mitigates catastrophic forgetting. Furthermore, knowledge distillation is applied during replay to reinforce the model’s ability to retain past capabilities while adapting to new data. To handle the diversity of mobility data, MoveGCL adopts a Mixture-of-Experts (MoE) architecture equipped with a mobility pattern-aware expert routing mechanism. This design enables the model to dynamically select expert modules tailored to local mobility characteristics. Together, these innovations make MoveGCL a practical and privacy-preserving solution for collaboratively building generalizable mobility foundation models across distributed, heterogeneous, and privacy-sensitive data sources.
In summary, our key contributions are as follows:
@@ 2.1. Mobility Data
Mobility data includes aggregated flows and individual trajectories (Barbosa et al., 2018; Yuan et al., 2025a; Zhang et al., 2025). Aggregated flows are relatively easier to obtain and have been widely used in urban analytics (Rong et al., 2024; Zhang et al., 2017). However, individual-level mobility data remain fragmented due to privacy concerns and institutional data silos (Kong et al., 2023; Yuan et al., 2025a). Real-world trajectory datasets, such as GeoLife (Zheng et al., [n. d.]), T-Drive (Yuan et al., 2010), NYC Taxi (New York City Taxi and Limousine Commission, 2023), and Foursquare (Yang et al., 2019), often have limited city coverage, short time spans, and sparse sampling. Some global-scale open datasets, such as the one used in UniTraj, have been introduced, but they suffer from low spatial-temporal resolution and inconsistent data quality. With the advancement of generative AI, synthetic mobility datasets have emerged, such as SynMob (Zhu et al., 2023), YJMob100K (Yabe et al., 2024) and WorldMove (Yuan et al., 2025b). However, the quality of synthetic data still falls short compared to real-world trajectories, particularly in terms of behavioral diversity, temporal continuity, and semantic consistency. In practice, access to real trajectory datasets typically requires signing NDAs, and most published studies do not release the mobility datasets they use due to privacy and legal restrictions (Schläpfer et al., 2021).
@@ 2.2. Mobility Foundation Models
Inspired by the success of foundation models in NLP and CV, recent efforts have explored pre-trained models for urban and mobility domains (Yuan et al., 2024d; Chai et al., 2025; Han et al., 2025; Zhou et al., 2024; Zhang et al., 2024; Choudhury et al., 2024). Early studies primarily focused on aggregated mobility data, leveraging mobility flows across cities to build unified spatio-temporal representations, and have demonstrated strong zero-shot transfer capabilities (Li et al., 2024b; Yuan et al., 2024b, a; Li et al., 2024a). In contrast, individual-level mobility foundation models are less developed. Researchers have explored multi-scale mobility modeling (Yuan et al., 2025a; Zhang et al., 2025; Long et al., 2024), aiming to capture both micro-level behaviors and macro-level patterns essential for generalization. Attempts such as UniTraj (Zhu et al., 2024), TrajBert (Si et al., 2023), and TrajFM (Lin et al., 2024) have explored learning from open trajectory datasets, but these datasets often consist of short-term or non-representative mobility traces that do not reflect regular human movement patterns. As a result, current models struggle to capture the full complexity and diversity of real-world individual mobility. Recently, LLMs have also widely utilized in generating human mobility (Shao et al., 2024; Gong et al., 2024; Ju et al., 2025; Feng et al., 2024; JIAWEI et al., 2024), but the gap between natural language and trajectory data suggests that mobility still requires native foundation models, which can later be aligned with LLMs to bridge symbolic reasoning and physical behavior modeling.
@@ 2.3. Continual Learning
Continual learning (Kim et al., 2022; Wang et al., 2024, 2024), also known as lifelong learning, aims to enable models to learn from a sequence of tasks or data streams without forgetting previously acquired knowledge (Yoon et al., 2021). A central challenge in continual learning is catastrophic forgetting (Li et al., 2019; Wickramasinghe et al., 2023), where the model’s performance on earlier tasks degrades as it learns new ones. Continual learning methods are typically categorized into three main types. Regularization-based methods introduce constraints on parameter updates to preserve important knowledge from earlier tasks, as seen in approaches like Elastic Weight Consolidation (EWC) and Synaptic Intelligence (SI). Replay-based methods mitigate forgetting by either storing a subset of previous data (experience replay) or generating pseudo-data (generative replay) to simulate past learning. Parameter isolation methods, on the other hand, allocate different parts of the model to different tasks, using techniques such as dynamic networks or task-specific masking to reduce interference between tasks.
@@ 4.2. Model Architecture
The Mixture-of-Experts (MoE) architecture comprises a router network and multiple expert networks, serving as a replacement for the Feed-Forward Network (FFN) within the Transformer (Masoudnia and Ebrahimpour, 2014). The output of the MoE layer, $F_{\mathrm{MoE}}(x)$, is the weighted sum of the selected expert outputs, where the weights are given by the router network’s output:
@@
The next-location prediction is performed by computing the similarity between the final output of the Mixture-of-Experts (MoE) Transformer blocks and the representation vectors of all candidate locations in the city. These location representations are generated using a Deep & Cross Network (DCN) (Wang et al., 2017), which consists of both a Cross layer and a Deep layer to capture feature interactions and nonlinear transformations.
@@ 5.1. Experimental Settings
  • Traditional approach: This includes Markov models (Gambs et al., 2012) that fit separate transition matrices for different datasets.
  • Deep mobility models: We include LSTM (Kong and Wu, 2018), Transformer (Vaswani et al., 2017), DeepMove (Feng et al., 2022), TrajBert (Si et al., 2023), and CLET (Lin et al., 2021) as representative baselines. For each dataset, we train a separate model to ensure fair comparison under the same training conditions.
  • Fedarated learning models: We evaluate against PMF (Feng et al., 2020) and LightTR (Liu et al., 2024a), which leverage federated learning frameworks for human mobility prediction while maintaining data decentralization. We apply their federated learning methods to train our model.
@@
Evaluation Metrics.
We adopt top-$k$ accuracy as our evaluation metric, defined as
@@ 5.2. Overall Performance
  • MoveGCL surpasses privacy-preserving federated learning approaches in both accuracy and stability. Compared to domain-specific baselines such as PVM (Feng et al., 2020) and LightTR (Liu et al., 2024a), MoveGCL achieves significantly higher performance. This advantage stems from its unified generative continual learning framework, which maintains global generalization without suffering from the synchronization overhead and convergence instability inherent in federated setups. This further supports its practicality in real-world multi-party mobility modeling scenarios.
@@ 5.4. Privacy Evaluation
As MoveGCL is built on a generative continual learning framework that does not retain raw data from previous cities, a key question is whether the synthetic data used for replay may inadvertently leak private information from the original training data. To rigorously evaluate the privacy-preserving properties of our approach, following the methodology in (Yuan et al., 2025a, 2024c), we conduct a comprehensive analysis from three complementary perspectives:
  • Uniqueness Testing (de Montjoye et al., 2013; Xu et al., 2017): To evaluate the degree of similarity between the generated data and the real data.
  • Membership Inference Attack (Lin et al., 2020a; Shokri et al., 2017): Given a trained model and a set of samples, it assesses whether an classifier can accurately determine which samples were included in the model’s training set based on the model’s outputs.
  • Differential Privacy (Abadi et al., 2016; Andrew et al., 2019): To ensure that the model does not depend on a small subset of training examples, we remove a minimal set of training samples and evaluate whether the distribution of model outputs undergoes an obvious change.
@@
Following the experimental setup in (Lin et al., 2020b; Shokri et al., 2017), we use the similarity between generated trajectories and their corresponding real input trajectories as the classification feature. For each trajectory $x$, MoveGCL generates $\tilde{x}$ autoregressively conditioned on $x$, and we compute a similarity score s(x,x~)s($\tilde{x}$) to form the input to the classifier. The classifier is then tasked with determining whether $x$ was included in the model’s training set. Positive samples consist of real‐world trajectories that were used during training, while negative samples are real trajectories from the same city that were held out. We evaluate the attack success rate, defined as the proportion of samples for which the classifier correctly infers membership status. We employ three widely used classification algorithms: Logistic Regression (LR), Support Vector Machine (SVM) and Random Forest (RF).
@@
In our experiment, we randomly select a subset of trajectories and consider two training scenarios: one in which this subset is included in the training data ($D$), and one in which it is excluded ($D^{\prime}$). For each scenario, we train $M$ on the corresponding dataset and then use each selected trajectory as a conditioning input to generate multiple synthetic trajectories. We compute a similarity score between each generated trajectory and its original conditioning trajectory. These similarity scores are modeled as two Gaussian distributions corresponding to $\mathbb{P}\bigl{[}M(z;\,D)\bigr{]}$ and $\mathbb{P}\bigl{[}M(z;\,D^{\prime})\bigr{]}$, respectively. Finally, we estimate the privacy‐budget parameters $\varepsilon$ from these distributions. As shown in Table 3, without applying any additional privacy-preserving mechanisms, MoveGCL naturally achieves a privacy budget of $\varepsilon\approx 1\text{--}2$ for 75% of randomly sampled trajectories. This level is generally considered an acceptable operating point for generative models (Lin et al., 2020a); for example, Apple adopts a privacy budget of $\varepsilon=4.0$ ***https://www.apple.com/privacy/docs/Differential_Privacy_Overview.pdf.

### v5

@@ Abstract.
Human mobility, foundation models, continual learning
@@ 1. Introduction
In natural language processing (NLP) (kojima2022large; brown2020language; bi2024deepseek) and computer vision (CV) (esser2024scaling; liu2024sora), the rise of large pre-trained foundation models (e.g., GPT, Sora) (radford2021learning) has significantly advanced model sharing and general-purpose intelligence through centralized training and open release. However, in the critical domain of human mobility modeling (barbosa2018human; luca2021survey), the era of large foundation models has not yet arrived. This is primarily due to the highly sensitive nature of mobility trajectory data, which involves personal privacy (kong2023mobility; yuan2025learning; yuan2022activity). Such data cannot be easily shared or jointly trained across institutions due to privacy constraints and legal restrictions. As a result, mobility datasets often exist in isolated silos (yuan2025worldmove; zhu2023synmob), with different organizations and researchers relying on their own private datasets, hindering data integration, benchmarking, and collaborative model development.
@@
To address the challenge of data, recent studies have proposed different training strategies to leverage multiple datasets, and Figure 1 compares different approaches. The most common strategy is to train models separately, as shown in Figure 1(a). Recent efforts such as UniTraj (zhu2024unitraj), TrajBert (si2023trajbert), and TrajFM (lin2024trajfm) have explored joint training for unified representation and cross-city generalization (Figure 1(b)), but these models remain tightly coupled with proprietary or limited-quality datasets. TrajFM and TrajBert are typically pre-trained on restricted or private data. While UniTraj uses public data, it suffers from low semantic richness and high sparsity. Federated learning (feng2020pmf; 10598171LightTR) offers a potential solution for distributed mobility model training (Figure 1(c)), but its reliance on frequent synchronization and communication poses challenges for scalability and practical deployment. Consequently, these approaches fall short of meeting the diverse, dynamic, and multi-source demands of real-world mobility modeling, and cannot support an open ecosystem of shared models as seen in NLP and CV domains.
@@
We propose MoveGCL, a scalable training framework for mobility foundation models based on generative continual learning. MoveGCL allows each data holder to evolve a shared model locally without exposing raw data, thereby ensuring full privacy preservation. Specifically, MoveGCL starts from a pre-trained base model and employs a synthetic trajectory replay mechanism: instead of accessing historical data, each participant generates synthetic trajectories that approximate previously seen mobility patterns. This replay process preserves prior knowledge and mitigates catastrophic forgetting. Furthermore, knowledge distillation is applied during replay to reinforce the model’s ability to retain past capabilities while adapting to new data. To handle the diversity of mobility data, MoveGCL adopts a Mixture-of-Experts (MoE) architecture equipped with a mobility pattern-aware expert routing mechanism. This design enables the model to dynamically select expert modules tailored to local mobility characteristics. Together, these innovations make MoveGCL a practical and privacy-preserving solution for collaboratively building generalizable mobility foundation models across distributed, heterogeneous, and privacy-sensitive data sources. In summary, our key contributions are as follows:
@@ 2.1. Mobility Data
Mobility data includes aggregated flows and individual trajectories (barbosa2018human; yuan2025learning; zhang2025noise). Aggregated flows are relatively easier to obtain and have been widely used in urban analytics (rong2024interdisciplinary; zhang2017deep). However, individual-level mobility data remain fragmented due to privacy concerns and institutional data silos (kong2023mobility; yuan2025learning). Real-world trajectory datasets, such as GeoLife (zheng2011geolife), T-Drive (yuan2010t), NYC Taxi (nyc_tlc_trip_data), and Foursquare (yang2019revisiting), often have limited city coverage, short time spans, and sparse sampling. Some global-scale open datasets, such as the one used in UniTraj, have been introduced, but they suffer from low spatial-temporal resolution and inconsistent data quality. With the advancement of generative AI, synthetic mobility datasets have emerged, such as SynMob (zhu2023synmob), YJMob100K (yabe2024yjmob100k) and WorldMove (yuan2025worldmove). However, the quality of synthetic data still falls short compared to real-world trajectories, particularly in terms of behavioral diversity, temporal continuity, and semantic consistency. In practice, access to real trajectory datasets typically requires signing NDAs, and most published studies do not release the mobility datasets they use due to privacy and legal restrictions (schlapfer2021universal).
@@ 2.2. Mobility Foundation Models
Inspired by the success of foundation models in NLP and CV, recent efforts have explored pre-trained models for urban and mobility domains (yuan2024urbandit; chai2025mobiworld; han2025trajmoe; zhou2024urban; zhang2024urban; choudhury2024towards). Early studies primarily focused on aggregated mobility data, leveraging mobility flows across cities to build unified spatio-temporal representations, and have demonstrated strong zero-shot transfer capabilities (li2024urbangpt; yuan2024uniflow; yuan2024unist; li2024opencity). In contrast, individual-level mobility foundation models are less developed. Researchers have explored multi-scale mobility modeling (yuan2025learning; zhang2025noise; long2024universal), aiming to capture both micro-level behaviors and macro-level patterns essential for generalization. Attempts such as UniTraj (zhu2024unitraj), TrajBert (si2023trajbert), and TrajFM (lin2024trajfm) have explored learning from open trajectory datasets, but these datasets often consist of short-term or non-representative mobility traces that do not reflect regular human movement patterns. As a result, current models struggle to capture the full complexity and diversity of real-world individual mobility. Recently, LLMs have also widely utilized in generating human mobility (shao2024beyond; gong2024mobility; ju2025trajllm; feng2024agentmove; jiawei2024large), but the gap between natural language and trajectory data suggests that mobility still requires native foundation models, which can later be aligned with LLMs to bridge symbolic reasoning and physical behavior modeling.
@@ 2.3. Continual Learning
Continual learning (kim2022theoretical; wang2024comprehensive; wang2024comprehensive), also known as lifelong learning, aims to enable models to learn from a sequence of tasks or data streams without forgetting previously acquired knowledge (yoon2021federated). A central challenge in continual learning is catastrophic forgetting (li2019learn; wickramasinghe2023continual), where the model’s performance on earlier tasks degrades as it learns new ones. Continual learning methods are typically categorized into three main types. Regularization-based methods introduce constraints on parameter updates to preserve important knowledge from earlier tasks, as seen in approaches like Elastic Weight Consolidation (EWC) and Synaptic Intelligence (SI). Replay-based methods mitigate forgetting by either storing a subset of previous data (experience replay) or generating pseudo-data (generative replay) to simulate past learning. Parameter isolation methods, on the other hand, allocate different parts of the model to different tasks, using techniques such as dynamic networks or task-specific masking to reduce interference between tasks.
@@ 4.2. Model Architecture
The Mixture-of-Experts (MoE) architecture comprises a router network and multiple expert networks, serving as a replacement for the Feed-Forward Network (FFN) within the Transformer (masoudnia2014mixture). The output of the MoE layer, $F_{\mathrm{MoE}}(x)$, is the weighted sum of the selected expert outputs, where the weights are given by the router network’s output:
@@
The next-location prediction is performed by computing the similarity between the final output of the Mixture-of-Experts (MoE) Transformer blocks and the representation vectors of all candidate locations in the city. These location representations are generated using a Deep & Cross Network (DCN) (wang2017deep), which consists of both a Cross layer and a Deep layer to capture feature interactions and nonlinear transformations.
@@ 5.1. Experimental Settings
  • Traditional approach: This includes Markov models (gambs2012next) that fit separate transition matrices for different datasets.
  • Deep mobility models: We include LSTM (kong2018hst), Transformer (vaswani2017attention), DeepMove (deepmove), TrajBert (si2023trajbert), and CLET (CLET) as representative baselines. For each dataset, we train a separate model to ensure fair comparison under the same training conditions.
  • Fedarated learning models: We evaluate against PMF (feng2020pmf) and LightTR (10598171LightTR), which leverage federated learning frameworks for human mobility prediction while maintaining data decentralization. We apply their federated learning methods to train our model.
@@ 5.2. Overall Performance
  • MoveGCL surpasses privacy-preserving federated learning approaches in both accuracy and stability. Compared to domain-specific baselines such as PVM (feng2020pmf) and LightTR (10598171LightTR), MoveGCL achieves significantly higher performance. This advantage stems from its unified generative continual learning framework, which maintains global generalization without suffering from the synchronization overhead and convergence instability inherent in federated setups. This further supports its practicality in real-world multi-party mobility modeling scenarios.
@@ 5.4. Privacy Evaluation
As MoveGCL is built on a generative continual learning framework that does not retain raw data from previous cities, a key question is whether the synthetic data used for replay may inadvertently leak private information from the original training data. To rigorously evaluate the privacy-preserving properties of our approach, following the methodology in (yuan2025learning; yuan2024needdyn), we conduct a comprehensive analysis from three complementary perspectives:
  • Uniqueness Testing (demontjoye2013unique; xu2017trajectory): To evaluate the degree of similarity between the generated data and the real data.
  • Membership Inference Attack (lin2020using; shokri2017membership): Given a trained model and a set of samples, it assesses whether an classifier can accurately determine which samples were included in the model’s training set based on the model’s outputs.
  • Differential Privacy (abadi2016deep; andrew2019tensorflow): To ensure that the model does not depend on a small subset of training examples, we remove a minimal set of training samples and evaluate whether the distribution of model outputs undergoes an obvious change.
@@
Following the experimental setup in (lin2020doppelganger; shokri2017membership), we use the similarity between generated trajectories and their corresponding real input trajectories as the classification feature. For each trajectory $x$, MoveGCL generates $\tilde{x}$ autoregressively conditioned on $x$, and we compute a similarity score s(x,x~)s($\tilde{x}$) to form the input to the classifier. The classifier is then tasked with determining whether $x$ was included in the model’s training set. Positive samples consist of real‐world trajectories that were used during training, while negative samples are real trajectories from the same city that were held out. We evaluate the attack success rate, defined as the proportion of samples for which the classifier correctly infers membership status. We employ three widely used classification algorithms: Logistic Regression (LR), Support Vector Machine (SVM) and Random Forest (RF).
@@
In our experiment, we randomly select a subset of trajectories and consider two training scenarios: one in which this subset is included in the training data ($D$), and one in which it is excluded ($D^{\prime}$). For each scenario, we train $M$ on the corresponding dataset and then use each selected trajectory as a conditioning input to generate multiple synthetic trajectories. We compute a similarity score between each generated trajectory and its original conditioning trajectory. These similarity scores are modeled as two Gaussian distributions corresponding to $\mathbb{P}\bigl[M(z;\,D)\bigr]$ and $\mathbb{P}\bigl[M(z;\,D^{\prime})\bigr]$, respectively. Finally, we estimate the privacy‐budget parameters $\varepsilon$ from these distributions. As shown in Table 3, without applying any additional privacy-preserving mechanisms, MoveGCL naturally achieves a privacy budget of $\varepsilon\approx 1\text{--}2$ for 75% of randomly sampled trajectories. This level is generally considered an acceptable operating point for generative models (lin2020using); for example, Apple adopts a privacy budget of $\varepsilon=4.0$ ***https://www.apple.com/privacy/docs/Differential_Privacy_Overview.pdf.
@@ Appendix A Dataset
Appendix B Evaluation Metrics
We adopt top-$k$ accuracy as our evaluation metric, defined as